# Improved Margin Generalization Bounds for Voting Classifiers

### v1

@@ 5 Final Upper bound
There exists a universal constant $c$ such that: For ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ a hypothesis class of VC-dimension $d$, distribution ${\mathcal{D}}$ over ${\mathcal{X}}\times\{-1,1\}$, failure parameter $10<\delta<$ 1, and training sequence size $m\in\mathbb{N}$, it holds with probability at least $1-\delta$ over ${\mathbf{S}}\sim{\mathcal{D}}^{m}$ that: for any margin $0<\gamma\leq 1$ and any function $f\in\operatorname{\Delta({\mathcal{H}})}$
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+\negmedspace\negmedspace\sqrt{c\cdot\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma^{2}}{d}\right)}}{m\gamma^{2}}+\frac{\ln{\left(\frac{e}{\delta}\right)}}{m}\right)}\negmedspace+\negmedspace c\negmedspace\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma^{2}}{d}\right)}}{m\gamma^{2}}+\frac{\ln{\left(\frac{e}{\delta}\right)}}{m}\right)\negmedspace.$
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Big{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{9600\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$ (41)
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{9600\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Big{)}.$
@@
We first notice that if $\frac{9600\tilde{c}^{2}\min(c,1)d}{32m}\geq\gamma^{2}$, the right-hand side of the above is greater than $1$, and we are done by the left-hand side always being at most $1$, i.e. the above inequality holds with probability $1$. Thus, we now consider the case where $1\geq\gamma^{2}\geq\frac{9600\tilde{c}^{2}\min(c,1)d}{32m}$. We now further restrict the values we have to consider for the bound in Eq. 41 to be non vacomuous, this is done to the end of employing Corollary 6 later in the argument.
@@
We will soon argue that the function $\operatorname{\Gamma}(x)/x$ is decreasing for $x\geq e^{4},$ we notice this is also a continuous function i $x.$ Now consider the function $9600\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}/(\frac{32m\gamma^{2}}{\min(c,1)d}),$ which by the above is decreasing in $\frac{32m\gamma^{2}}{\min(c,1)d}\geq e^{e},$ which is the case since we consider $\gamma$ such that $1\geq\frac{32m\gamma^{2}}{\min(c,1)d}\geq 9600\tilde{c}^{2}\geq 9600$ This also implies that for the right hand side of Eq. 41 being less than $1$ for any $1\geq\gamma^{2}\geq\frac{9600\tilde{c}^{2}\min(c,1)d}{32m}$ it must be the case that $m$ is such that $19600\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}/(\frac{32m}{\min(c,1)d})<$ 1 which we assume is the case from now on. Furthermore, by the above argued continuity and the function being decreasing, let $\gamma^{\prime}$ be such that $1\geq\gamma^{\prime 2}\geq\frac{9600\tilde{c}^{2}\min(c,1)d}{32m}$ and that $9600\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}/(\frac{32m\gamma^{2}}{\min(c,1)d})=1,$ notice that $\gamma^{\prime}$ must be within the interval since the function was decreasing in $\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)$ and for the value $\gamma^{2}=\frac{9600\tilde{c}^{2}\min(c,1)d}{32m}$ the function is greater than $1$ Notice that for the values of $\gamma,$ between $\gamma^{\prime 2}\geq\gamma\geq\frac{9600\tilde{c}^{2}\min(c,1)d}{32m},$ the bound on the right hand side of Eq. 41 is again larger than $1,$ so we consider from now on the case $1\geq\gamma^{2}\geq\gamma^{\prime 2},$ and $m$ such that $19600\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}/(\frac{32m}{\min(c,1)d})<$ 1 . Furthermore, again by the above argued monotonicity we notice that for $1\geq\gamma^{2}\geq\gamma^{\prime 2}$ we have that
@@
$1\displaystyle 9600\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}/\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)\leq 9600\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32\gamma^{\prime 2}m}{\min(c,1)d}\right)}/\left(\frac{32\gamma^{\prime 2}m}{\min(c,1)d}\right)=$ 1 (42)
@@
To argue that $\operatorname{\Gamma}(x)/x$ is decreasing for $x\geq e^{4},$ we notice that for $x\geq e^{4}$ we have that $\operatorname{\Gamma}(x)=\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}(x))\operatorname{\mathrm{Ln}}(x)=\ln^{2}(\ln{\left(x\right)})\ln{\left(x\right)},$ thus it suffices to show that the function $f(x)=\ln^{2}(\ln{\left(x\right)})\ln{\left(x\right)}/x$ is decreasing for $x\geq e^{4}.$ $f$ has derivative $(2\ln{\left(\ln{\left(x\right)}\right)}-\ln{\left(x\right)}(\ln{\left(\ln{\left(x\right)}\right)})^{2}+\left(\ln{\left(\ln{\left(x\right)}\right)}\right)^{2})/x^{2}.$ We observe that for $x\geq e^{2}$ we have that $-\ln{\left(x\right)}(\ln{\left(\ln{\left(x\right)}\right)})^{2}/2+\left(\ln{\left(\ln{\left(x\right)}\right)}\right)^{2}\leq 0$ and for $x\geq e^{4}$ we have that $2\ln{\left(\ln{\left(x\right)}\right)}-\ln{\left(x\right)}(\ln{\left(\ln{\left(x\right)}\right)})^{2}/2\leq 0$ Whereby we conclude that $f(x)=\ln^{2}(\ln{\left(x\right)})\ln{\left(x\right)}/x$ is decreasing for $x\geq e^{4}.$ We are now ready to setup the parameters for the union bound over the different level sets of $\gamma$
Let $I=\{0,\ldots,\lfloor\log_{2}(1/\gamma^{\prime})\rfloor\}$ and define $\gamma_{0}^{i}=\sqrt{2^{i}\gamma^{\prime}}$ and $\gamma_{1}^{i}=\sqrt{2}\gamma_{0}^{i}$, for $i\in I$, except $i=\lfloor\log_{2}(1/\gamma^{\prime})\rfloor,$ where we define $\gamma_{1}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}=1$. For $i\in I$ we also define $N_{i}=\exp{\left(\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}$ and $\delta_{i}=\frac{\delta}{1640N_{i}^{2}}\frac{d}{m(\gamma_{0}^{i})^{2}}$. Furthermore, for each $i\in I$ we define $J_{i}=\{0,1,\ldots,\big{\lfloor}\frac{m}{\ln{\left(N_{i}\right)}}\big{\rfloor}\}$ and for $j\in J_{i}$ define $\tau_{0}^{i,j}=j\frac{\ln{\left(N_{i}\right)}}{m}$ and $\tau_{1}^{i,j}=\left(j+1\right)\frac{\ln{\left(N_{i}\right)}}{m}$. Lastly for $i\in I$ and $j\in J_{i}$ we define
@@
$\leq\left(\left\lfloor\frac{m}{\ln{\left(N_{i}\right)}}\right\rfloor+1\right)\delta_{i}\sup_{X\in{\mathcal{X}}^{2m}}{\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2}).$ (43)
We remark that if $i\in I$ is such that $\frac{\ln{\left(N_{i}\right)}}{m}>1$, then we have that $\tau_{1}^{i,j}=(j+1)\frac{\ln{\left(N_{i}\right)}}{m}>1,$ which implies that we could have upper bounded the above with 00. Thus, we assume for now that $i\in I$ such that $\frac{\ln{\left(N_{i}\right)}}{m}\leq 1$, which implies that $\big{(}\big{\lfloor}\frac{m}{\ln{\left(N_{i}\right)}}\big{\rfloor}+1\big{)}$ is upper bounded by $\frac{2m}{\ln{\left(N_{i}\right)}}$.
@@
where the inequality follows from in the case that $\gamma=\min(2\gamma_{1}^{i},1)=1,$ then since the functions in $\operatorname{\Delta({\mathcal{H}})}$ only attains values in $[-1,1]$, we have that $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil}=\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 1\right\rceil}=\operatorname{\Delta({\mathcal{H}})}$ and that the covering number being decreasing in the precision argument and $\gamma_{0}^{i}/2\geq\alpha\gamma$. Furthermore, to the end of invoking Corollary 6, we now argue that $\gamma^{2}\geq\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}$ for the above $\gamma$ and $\alpha.$ To this end we notice that since $\gamma_{0}^{i}$ and $\gamma\sqrt{2}\gamma_{0}^{i},$ except $i=\lfloor\log_{2}(1/\gamma^{\prime})\rfloor,$ where we define $\gamma_{1}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}=1$, we have the following $\frac{1}{4\sqrt{2}}\leq\alpha=\frac{\gamma_{0}^{i}}{4\gamma_{1}^{i}}\leq\frac{1}{4}.$ Furthermore, we argued previously that it suffices to consider the case that $\frac{9600\tilde{c}^{2}\min(c,1)d}{32}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}/{m}<1.$ Now if $\gamma=1$ we get that
$\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}\leq\frac{128cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{2m}{cd})})}{m}\leq\frac{9600\tilde{c}^{2}\min(c,1)d}{32m}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}<1=\gamma^{2}$
where the first inequality uses that $\frac{1}{4\sqrt{2}}\leq\alpha\leq\frac{1}{4},$ and the second inequality that $c\leq\tilde{c}=\max(c^{\prime},1/c),$ where $c^{\prime}\geq c$ and $9600/32\geq 128$ and the last inequality the case $\gamma=1$ we consider.
@@
$\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}\leq\frac{128cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8m(\gamma_{1}^{i})^{2}}{cd})})}{m}$ $\leq\frac{9600\tilde{c}^{2}\min(c,1)d}{32m}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{1}^{i})^{2}}{\min(c,1)d}\right)}$
@@
where the first inequality uses that $\frac{1}{4\sqrt{2}}\leq\alpha\leq\frac{1}{4},$ and the second inequality that $c\leq\tilde{c}=\max(c^{\prime},1/c),$ where $c^{\prime}\geq c,$ that $\gamma=2\gamma_{1}^{i}$ and $9600/32\geq 128$ where the second to last inequality follows from Eq. 42 and $1\geq\gamma_{1}^{i}\geq\gamma_{0}^{i}\geq\gamma^{\prime}$ and the last inequality by $\gamma=2\gamma_{1}^{i}$, whereby we have argued that $\gamma^{2}\geq\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}$, and since $0<\gamma\leq 1$ and $20<\alpha<1/$ 2 we get by invoking Corollary 6 that
@@
${\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2})\leq{\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil},\alpha\gamma)$ $\leq\exp{\left(\frac{c^{\prime}d}{\alpha^{2}\gamma^{2}}\operatorname{\Gamma}\left(\frac{8\alpha m\gamma^{2}}{cd}\right)\right)}$
$\leq\exp{\left(\frac{16c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}\negmedspace{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}\negmedspace\right)}\leq N_{i},$
@@
where the first inequality follows from Eq. 44, the second inequality by Corollary 6, the third inequality by using on the term outside of $\operatorname{\Gamma}(\cdot)$ that $\frac{1}{\alpha^{2}\gamma^{2}}=\frac{16(\gamma_{1}^{i})^{2}}{(\gamma_{0}^{i})^{2}\min(4(\gamma_{1}^{i})^{2},1)}\leq 16/(\gamma_{0}^{i})^{2}$ and using on the term inside of $\operatorname{\Gamma}(\cdot)$ that $\gamma_{1}^{i}\leq\sqrt{2}(\gamma_{0}^{i})^{2}$ to upper bound $\alpha\gamma^{2}=\frac{\gamma_{0}^{i}}{4\gamma_{1}^{i}}(\min(2\gamma_{1}^{i},1))^{2}\leq\sqrt{2}(\gamma_{0}^{i})^{2}$. Thus, by using that we concluded that ${\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2})\leq N_{i}$ and that $(\lfloor\frac{m}{\ln{(N_{i})}}\rfloor+1)\leq\frac{2m}{\ln{(N_{i})}}$, and $\delta_{i}=\frac{\delta}{1640N_{i}^{2}}\frac{d}{m(\gamma_{0}^{i})^{2}},$ we get by plugging into the last expression of Section 5, that

### v2

@@ 5 Final Upper bound
There exists a universal constant $C>0$ such that: For ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ a hypothesis class of VC-dimension $d$, distribution ${\mathcal{D}}$ over ${\mathcal{X}}\times\{-1,1\}$, failure parameter $10<\delta<$ 1, and training sequence size $m\in\mathbb{N}$, it holds with probability at least $1-\delta$ over ${\mathbf{S}}\sim{\mathcal{D}}^{m}$ that: for any margin $0<\gamma\leq 1$ and any function $f\in\operatorname{\Delta({\mathcal{H}})}$
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+\sqrt{C\cdot\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma^{2}}{d}\right)}}{m\gamma^{2}}+\frac{\ln{\left(\frac{e}{\delta}\right)}}{m}\right)}+C\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma^{2}}{d}\right)}}{m\gamma^{2}}+\frac{\ln{\left(\frac{e}{\delta}\right)}}{m}\right).$
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Big{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{20000\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$ (41)
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{20000\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Big{)}.$
@@
We first notice that if $\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}\geq\gamma^{2}$, the right-hand side of the above is greater than $1$, and we are done by the left-hand side always being at most $1$, i.e. the above inequality holds with probability $1$. Thus, we now consider the case where $1\geq\gamma^{2}\geq\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}$. We now further restrict the values we have to consider for the bound in Eq. 41 to be non vacomuous, this is done to the end of employing Corollary 6 later in the argument.
@@
We will soon argue that the function $\operatorname{\Gamma}(x)/x$ is strictly decreasing for $x\geq e^{4},$ furthermore it is also a continuous function for $x>0.$ We now consider the function $20000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}/(\frac{32m\gamma^{2}}{\min(c,1)d}),$ which is strictly decreasing in $\frac{32m\gamma^{2}}{\min(c,1)d}$ when $\frac{32m\gamma^{2}}{\min(c,1)d}\geq e^{4},$ which is the case since we consider $\gamma$ such that $\frac{32m\gamma^{2}}{\min(c,1)d}\geq 20000\tilde{c}^{2}\geq 20000.$ This also implies that for the right hand side of Eq. 41 to be less than $1$ for any $1\geq\gamma^{2}\geq\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}$ it must be the case that $m$ is such that $120000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}/(\frac{32m}{\min(c,1)d})<$ 1 which we assume is the case from now on. Now let $\gamma^{\prime}$ be such that $1\geq\gamma^{\prime 2}\geq\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}$ and that $120000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{\prime 2}}{\min(c,1)d}\right)}/(\frac{32m\gamma^{\prime 2}}{\min(c,1)d})=$ 1 . We notice that $\gamma^{\prime}$ exists uniquely in this interval since the function $20000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}/(\frac{32m\gamma^{2}}{\min(c,1)d})$ was strictly decreasing(and continuous) in $\frac{32m\gamma^{2}}{\min(c,1)d}$ for $\frac{32m\gamma^{2}}{\min(c,1)d}\geq e^{4}$ and for the value $\gamma^{2}=\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}$, we have that $\frac{32m\gamma^{2}}{\min(c,1)d}\geq e^{4}$ and the function evaluated at $\gamma^{2}=\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}$ is greater than $1$ and for $\gamma=1$, the function evalutes to something strictly below $1$ Notice that for the values of $\gamma,$ between $\gamma^{\prime 2}\geq\gamma^{2}\geq\frac{20000\tilde{c}^{2}\min(c,1)d}{32m},$ the bound on the right hand side of Eq. 41 is again larger than $1,$ by the $20000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}/(\frac{32m\gamma^{2}}{\min(c,1)d})$ being strictly decreasing for these values, so we consider from now on the case $1\geq\gamma^{2}\geq\gamma^{\prime 2},$ and $m$ such that $120000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}/(\frac{32m}{\min(c,1)d})<$ 1 . Furthermore, again by the above argued monotonicity we notice that for $1\geq\gamma^{2}\geq\gamma^{\prime 2}$ we have that
@@
$1\displaystyle 20000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}/\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)\leq 20000\tilde{c}^{2}\operatorname{\Gamma}{\left(\frac{32\gamma^{\prime 2}m}{\min(c,1)d}\right)}/\left(\frac{32\gamma^{\prime 2}m}{\min(c,1)d}\right)=$ 1 (42)
@@
To argue that $\operatorname{\Gamma}(x)/x$ is strictly decreasing for $x\geq e^{4},$ we notice that for $x\geq e^{4}$ we have that $\operatorname{\Gamma}(x)=\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}(x))\operatorname{\mathrm{Ln}}(x)=\ln^{2}(\ln{\left(x\right)})\ln{\left(x\right)},$ thus it suffices to show that the function $f(x)=\ln^{2}(\ln{\left(x\right)})\ln{\left(x\right)}/x$ is strictly decreasing for $x\geq e^{4}.$ $f$ has derivative $(2\ln{\left(\ln{\left(x\right)}\right)}-\ln{\left(x\right)}(\ln{\left(\ln{\left(x\right)}\right)})^{2}+\left(\ln{\left(\ln{\left(x\right)}\right)}\right)^{2})/x^{2}.$ We observe that for $x\geq e^{4}$ we have that $-\ln{\left(x\right)}(\ln{\left(\ln{\left(x\right)}\right)})^{2}/2+\left(\ln{\left(\ln{\left(x\right)}\right)}\right)^{2}<0$ and for $x\geq e^{4}$ we have that $2\ln{\left(\ln{\left(x\right)}\right)}-\ln{\left(x\right)}(\ln{\left(\ln{\left(x\right)}\right)})^{2}/2\leq 0$ Whereby we conclude that the derivative of $f(x)=\ln^{2}(\ln{\left(x\right)})\ln{\left(x\right)}/x$ is negative for $x\geq e^{4}$ so strictly decreasing for $x\geq e^{4}.$ We are now ready to setup the parameters for the union bound over the different level sets of $\gamma$
Let $I=\{0,\ldots,\lfloor\log_{2}(1/\gamma^{\prime})\rfloor\}$ and define $\gamma_{0}^{i}=2^{i}\gamma^{\prime}$ and $\gamma_{1}^{i}=2\gamma_{0}^{i}$, for $i\in I$, except $i=\lfloor\log_{2}(1/\gamma^{\prime})\rfloor,$ where we define $\gamma_{1}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}=1$. For $i\in I$ we also define $N_{i}=\exp{\left(\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}$ and $\delta_{i}=\frac{\delta}{1640N_{i}^{2}}\frac{d}{m(\gamma_{0}^{i})^{2}}$. Furthermore, for each $i\in I$ we define $J_{i}=\{0,1,\ldots,\big{\lfloor}\frac{m}{\ln{\left(N_{i}\right)}}\big{\rfloor}\}$ and for $j\in J_{i}$ define $\tau_{0}^{i,j}=j\frac{\ln{\left(N_{i}\right)}}{m}$ and $\tau_{1}^{i,j}=\left(j+1\right)\frac{\ln{\left(N_{i}\right)}}{m}$. Lastly for $i\in I$ and $j\in J_{i}$ we define
@@
$\leq\left(\left\lfloor\frac{m}{\ln{\left(N_{i}\right)}}\right\rfloor+1\right)\delta_{i}\sup_{X\in{\mathcal{X}}^{2m}}{\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2}).$
We remark that if $i\in I$ is such that $\frac{\ln{\left(N_{i}\right)}}{m}\geq 1$, then we have that $\tau_{1}^{i,j}=(j+1)\frac{\ln{\left(N_{i}\right)}}{m}\geq 1,$ which implies, by $\beta_{i,j}>0$, and $\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq 1$, that we could have upper bounded the above probability with 00. Thus, in this case any non-negative number is an upper bound of the above. If $i\in I$ is such that $\frac{\ln{\left(N_{i}\right)}}{m}<1$, we have that $\big{(}\big{\lfloor}\frac{m}{\ln{\left(N_{i}\right)}}\big{\rfloor}+1\big{)}$ is upper bounded by $\frac{2m}{\ln{\left(N_{i}\right)}}$. Thus, we conclude that
$\leq\frac{2m}{\ln{\left(N_{i}\right)}}\delta_{i}\sup_{X\in{\mathcal{X}}^{2m}}{\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2}).$ (43)
@@
where the inequality in the case that $\gamma=\min(2\gamma_{1}^{i},1)=1,$ follows from the functions in $\operatorname{\Delta({\mathcal{H}})}$ only attaining values in $[-1,1]$, so $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil}=\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 1\right\rceil}=\operatorname{\Delta({\mathcal{H}})}$ and the covering number being decreasing in the precision argument and $\gamma_{0}^{i}/2\geq\alpha\gamma$ when $\gamma=\min(2\gamma_{1}^{i},1)=1$, in the case that $\gamma=2\gamma_{1}^{i}$ the above holds with equality. Furthermore, to the end of invoking Corollary 6(with $2m$ points), we now argue that $\gamma^{2}\geq\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha(2m)\gamma^{2}}{cd})})}{\alpha^{2}(2m)}$ for the above $\gamma$ and $\alpha.$ To this end we notice that since $\gamma_{0}^{i}$ and $\gamma_{1}^{i}=2\gamma_{0}^{i},$ except $i=\lfloor\log_{2}(1/\gamma^{\prime})\rfloor,$ where we define $\gamma_{1}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}=1$, we have that $\frac{1}{8}\leq\alpha=\frac{\gamma_{0}^{i}}{4\gamma_{1}^{i}}\leq\frac{1}{4}.$ Furthermore, we argued previously that it suffices to consider the case that $\frac{20000\tilde{c}^{2}\min(c,1)d}{32}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}/{m}<1.$ Now if $\gamma=1$ we get that
$\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha(2m)\gamma^{2}}{cd})})}{\alpha^{2}(2m)}\leq\frac{128cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{4m}{cd})})}{m}\leq\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}\operatorname{\Gamma}{\left(\frac{32m}{\min(c,1)d}\right)}<1=\gamma^{2}$
where we in the first inequality use that $\frac{1}{8}\leq\alpha\leq\frac{1}{4},$ and in the second inequality use that $c\leq\tilde{c}=\max(c^{\prime},1/c),$ where $c^{\prime}\geq c$ and $c^{\prime}\geq 1$, and $20000/32\geq 128$ and we use in the last inequality that we consider the case $\gamma=1$.
@@
$\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha(2m)\gamma^{2}}{cd})})}{\alpha^{2}(2m)}\leq\frac{128cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{16m(\gamma_{1}^{i})^{2}}{cd})})}{m}$ $\leq\frac{20000\tilde{c}^{2}\min(c,1)d}{32m}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{1}^{i})^{2}}{\min(c,1)d}\right)}$
@@
where in the first inequality we use that $\frac{1}{8}\leq\alpha\leq\frac{1}{4},$ and that $\gamma=2\gamma_{1}^{i}$, and in the second inequality we use that $c\leq\tilde{c}=\max(c^{\prime},1/c),$ where $c^{\prime}\geq c$ and $c^{\prime}\geq 1$, and $20000/32\geq 128$, furthermore the second to last inequality follows from Eq. 42 and $1\geq\gamma_{1}^{i}\geq\gamma_{0}^{i}\geq\gamma^{\prime}$ and the last inequality follows by $\gamma=2\gamma_{1}^{i}$, whereby we have argued that $\gamma^{2}\geq\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha(2m)\gamma^{2}}{cd})})}{\alpha^{2}(2m)}$, and since $0<\gamma\leq 1$ and $20<\alpha<1/$ 2 we get by invoking Corollary 6(with $2m$ points) that
@@
${\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2})\leq{\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil},\alpha\gamma)$ $\leq\exp{\left(\frac{c^{\prime}d}{\alpha^{2}\gamma^{2}}\operatorname{\Gamma}\left(\frac{8\alpha(2m)\gamma^{2}}{cd}\right)\right)}$
$\leq\exp{\left(\frac{16c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}\leq N_{i},$
@@
where the first inequality follows from Eq. 44, the second inequality by Corollary 6, the third inequality by using on the term outside of $\operatorname{\Gamma}(\cdot)$ that $\frac{1}{\alpha^{2}\gamma^{2}}=\frac{16(\gamma_{1}^{i})^{2}}{(\gamma_{0}^{i})^{2}\min(4(\gamma_{1}^{i})^{2},1)}\leq 16/(\gamma_{0}^{i})^{2}$ and using on the term inside of $\operatorname{\Gamma}(\cdot)$ that $\gamma_{1}^{i}\leq 2\gamma_{0}^{i}$ to upper bound $\alpha\gamma^{2}=\frac{\gamma_{0}^{i}}{4\gamma_{1}^{i}}(\min(2\gamma_{1}^{i},1))^{2}\leq 2(\gamma_{0}^{i})^{2}$. Thus, by using that we concluded that ${\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2})\leq N_{i},$ and that $\delta_{i}=\frac{\delta}{1640N_{i}^{2}}\frac{d}{m(\gamma_{0}^{i})^{2}},$ we get by plugging into the last expression of Section 5, that